# OccMamba: Semantic Occupancy Prediction with State Space Models

### v1

@@ 3.3 Hierarchical Mamba Module with Height-prioritized Reordering
Semantic occupancy prediction is challenging due to the high dimensionality and density of voxel grids, often involving millions of voxels. Previous methods (Huang et al. 2023; Ma et al. 2024) resort to projection techniques but suffer from information loss. In contrast, Mamba’s linear computational complexity enables direct processing of large voxel features, avoiding the limitations of deformable attention and convolutional methods, such as the need for key point selection or limited receptive fields.
@@
Height-prioritized Reordering Scheme. Before feeding 3D voxel features into the Mamba block, it’s necessary to reorder them into 1D sequences. A poor reordering strategy can disrupt the intrinsic spatial relationships between adjacent voxels in the 3D space, especially when dealing with a large number of voxel grids, thereby impacting Mamba’s performance. Inspired by the Hilbert curve (Hilbert 2013), we propose a height-prioritized 2D Hilbert expansion, tailored to the flat spatial structure typical of semantic occupancy prediction tasks, where height information provides valuable categorical clues, as it often correlates with object categories, reveals terrain features, and distinguishes different spatial regions in the environment. Specifically, we divide the ”xyz” coordinates into the ”xy” plane and a ”z” dimension. Starting on the xy-plane at z=0, the process extends vertically along the z-axis, creating vertical sequences. These sequences are then ordered and interconnected following the 2D Hilbert curve on the xy plane, forming the height-prioritized 2D Hilbert expansion, as illustrated in Fig. 3(d). For comparison, Fig. 3(a)-(c) illustrate the results of reordering according to the ’xyz’ sequence, the ’zxy’ sequence, and the 3D Hilbert curve, respectively. Our reordering strategy prioritizes z-axis spatial information while maintaining strong spatial proximity in the xy plane. By this way, we can better leverage Mamba’s contextual modeling.
@@
Hierarchical Mamba Encoder and Decoder. To effectively utilize the contextual information across spatial resolutions, we introduce the hierarchical Mamba encoder and decoder. The encoder comprises multiple groups, each with two Mamba blocks and a downsampling operation between groups for multi-scale voxel features. Correspondingly, the Mamba decoder mirrors this structure, utilizing paired Mamba blocks and upsampling operations. The upsampling operation incorporates skip connection from the corresponding layer in the Mamba encoder, ensuring consistency in feature dimensions across scales. Crucially, the reordering scheme is applied both before and after each Mamba blocks group, maintaining voxel representation during both downsampling and upsampling phases. As to the Mamba block, it consists of LayerNorm(LN), linear layers, 1D convolution, Silu activation, Selective SSM, and residual connections. Given an input $\mathbf{V}\in\mathbb{R}^{\mathcal{B}\times\mathcal{N}\times\mathcal{C}}$, where $\mathcal{B}$, $\mathcal{N}$ and $\mathcal{C}$ denote the batch size, the number of 1D sequences, and the Mamba feature dimension, respectively, the output $\mathbf{V^{\prime}}$ can be computed as below:
@@
$\mathbf{V^{\prime}}$ $=\text{Linear}(\mathbf{V^{\prime}_{2}}\odot\text{Silu}(\text{Linear}(\mathbf{V^{\prime}_{1}}))),$
where $\odot$ represents the dot product.
Endowed with the hierarchical Mamba encoder and decoder, our OccMamba utilizes features at multiple resolutions, thereby leveraging both coarse and fine-grained spatial information. This multi-scale approach preserves low-resolution details, enhancing contextual modeling and mitigating performance degradation with large data volumes.
@@ 3.4 Training Objective
Our ultimate training objective is composed of five terms, including the cross-entropy loss $\mathcal{L}_{\text{CE}}$, the lovasz-softmax loss $\mathcal{L}_{\text{iou}}$ (Berman, Triki, and Blaschko 2018), the geometric and semantic affinity loss $\mathcal{L}_{\text{geo}}$ and $\mathcal{L}_{\text{sem}}$ (Cao and De Charette 2022), and the depth supervision loss $\mathcal{L}_{\text{depth}}$ (Li et al. 2023a). These terms are crucial for optimizing our model’s performance in various aspects: $\mathcal{L}_{\text{CE}}$ ensures accurate classification, $\mathcal{L}_{\text{iou}}$ enhances semantic segmentation, $\mathcal{L}_{\text{geo}}$ improves spatial alignment, $\mathcal{L}_{\text{sem}}$ refines semantic understanding, and $\mathcal{L}_{\text{depth}}$ guides depth estimation and spatial relationships. The ultimate training objective is presented as follows:
@@ 4.1 Experimental Setup
Benchmarks. We conduct experiments on OpenOccupancy (Wang et al. 2023b), SemanticKITTI (Behley et al. 2019) and SemanticPOSS (Pan et al. 2020). OpenOccupancy is built upon the nuScenes dataset (Caesar et al. 2020), inheriting the data format of nuScenes. It comprises 700 training sequences and 150 validation sequences, with annotations for 17 classes. The occupancy annotations are represented in a 512$\times$512$\times$40 voxel grid, with each voxel sized at 0.2 meters. Notably, each frame in OpenOccupancy has a data range four times larger than the other datasets, imposing a heavier computational burden. In SemanticKITTI, sequences 00-10 (excluding 08), 08, and 11-21 are allocated for training, validation, and testing, respectively. For the occupancy annotations, a 256$\times$256$\times$32 grid is used, with voxels measuring 0.2 meters each. After pre-processing, a total of 19 classes are utilized for training and evaluation. SemanticPOSS, which closely resembles SemanticKITTI, employs sequences 00-05 and 02 as training and validation sets, respectively, with annotations for 11 classes.
Method Input Modality IoU mIoU
M-CONet C&L 26.8 20.5
OccMamba (ours) C 21.6 13.1
OccMamba (ours) L 36.8 23.1
OccMamba (ours) C&L 36.0 27.0
Table 2: Quantitative comparisons on OpenOccupancy validation set with v0.1 annotations. The best and second-best are in bold and underlined, respectively.
Method Input Modality mIoU
MonoScene C 11.1
SurroundOcc C 11.9
OccFormer C 12.3
RenderOcc C 12.8
LMSCNet L 17.0
JS3C-Net L 23.8
SSC-RS L 24.2
Co-Occ C&L 24.4
M-CONet C&L 20.4
OccMamba(ours) C&L 24.6
Method mIoU
SSCNet 15.2
LMSCNet 16.5
MotionSC 17.6
JS3C-Net 22.7
OccMamba(ours) 23.4
Implementation details. We use ResNet-50 (He et al. 2016) as the image backbone. Both the Mamba Encoder and Mamba Decoder consist of four groups and each group contains two Mamba blocks. For the OpenOccupancy dataset, we maintain identical setting to M-CONet (Wang et al. 2023b): for each frame, we utilize six surround view camera images as input, coupled with a fusion of ten frames of LiDAR points spanning the range of [-51.2m, 51.2m] along the X and Y axes, and [-2.0m, 6.0m] along the Z axis. Within the Mamba Module, we set the feature dimension to 384. On the SemanticKITTI dataset, our inputs consist of forward-facing stereo camera images alongside a single-frame LiDAR points, with their spatial extents defined as [0m, -25.6m, -2m, 51.2m, 25.6m, 4.4m]. We set the mamba feature dimension to 128 and employ Test Time Augmentation (TTA) to boost the performance. The used augmentations include the flipping of the X and Y axes, and four augmented inputs (including the original input) are sent to the model. Finally, for the SemanticPOSS dataset, we exclusively leverage single-frame LiDAR points like SemanticKITTI.
4.2 Experimental Results
OpenOccupancy. As shown in Table 1, we compare OccMamba on OpenOccupancy with previous methods under the settings of camera-only (Cao and De Charette 2022; Huang et al. 2023; Chen et al. 2020; Li et al. 2020a), LiDAR-only (Roldao, de Charette, and Verroust-Blondet 2020; Yan et al. 2021), and multi-modal fusion (Wang et al. 2023b; Pan, Wang, and Wang 2024). Our OccMamba achieves the best performance across all settings. Particularly, in the multi-modal fusion, OccMamba outperforms Co-Occ (Pan, Wang, and Wang 2024) by 3.1% in IoU and 3.2% in mIoU, respectively. These results underscore the efficacy of our approach in handling large-scale scenarios. Besides, we also conduct experiments on the newly published v0.1 annotations of OpenOccupancy in Table 2. OccMamba still exhibits competitive performance, achieving an mIoU of 27.0%, which is 6.5% higher than M-CONet (Wang et al. 2023b).
@@ 4.2 Experimental Results
SemanticKITTI & SemanticPOSS. As evident from Table 3, our OccMamba still outperforms the second-best method, i.e., Co-Occ (Pan, Wang, and Wang 2024), by 0.2% mIoU in SemanticKITTI test set. Note that we do not compare our OccMamba with SCPNet (Xia et al. 2023) and Occfiner (Shi et al. 2024) as they employ tricks such as label rectification, knowledge distillation or multi-frame concatenation. Strong performance is also observed in SemanticPOSS, as shown in Table 4. These results sufficiently demonstrate the superiority of our approach. The detailed classwise performance is put in the supplementary material.
Visual comparisons. From Fig. 4, our OccMamba produces more accurate predictions than M-CONet (Wang et al. 2023b) on OpenOccupancy validation set. Our OccMamba not only provides a more detailed representation of object shapes and semantics, but also performs preciser predictions on occluded objects and surfaces.
@@ 4.3 Ablation Study
The reported results are on the OpenOccupancy validation set with v0.1 annotations unless otherwise specified.
Reordering Schemes mIoU
XYZ sequence 25.3
3D Hilbert 25.5
ZXY sequence 25.5
Height-prioritized 2D Hilbert expansion 25.9
Table 5: Ablation study on different reordering schemes.
Reordering Schemes. To accelerate the training process, we set the mamba feature dimension to 128 and then compare our method with three alternative methods: ordering by XYZ sequence, ordering by ZXY sequence, and ordering using the 3D Hilbert curve. As shown in Table 5, our method achieves the highest mIoU of 25.9%. Additionally, it is apparent that reordering schemes that prioritize the z-axis outperform the other methods. This could be attributed to the fact that in semantic occupancy prediction, the overall shape of the scenes typically manifests as a flattened cuboid. Therefore, reordering schemes that prioritize the z-axis allow Mamba to better capture spatial proximity information, leading to improved prediction performance.
Method mIoU
MonoScene 11.1
MonoScene + Hierarchical Mamba module 11.7
@@
Generalization. We apply our hierarchical Mamba module to the MonoScene(Cao and De Charette 2022) framework. Specifically, we replace the 3D voxel decoder in this framework with our implementation and train on 25% of the SemanticKITTI training set. As shown in Table 6, our Mamba module can improve the performance of MonoScene from 11.1% mIoU to 11.7% mIoU on SemanticKITTI val set.
Memory Usage and Inference Time.
@@
M-CONet 37.3 12.1 416.0 20.5
OccMamba-384 29.9 8.5 398.1 27.0
OccMamba-128 19.2 6.2 235.7 25.9
Table 7: Comparison of memory usage and inference time. OccMamba-384 means OccMamba with mamba feature dimension being 384.
In Table 7, we present a comparative analysis of memory usage and inference time between OccMamba and M-CONet during both training and inference phases. The experiment is trained on 8 A40 GPUs and performed inference on a single RTX 4090 GPU. When the mamba feature dimension is set to 384, OccMamba demonstrates approximately a 20% reduction in memory usage during training and about a 30% reduction during inference compared to M-CONet. When the mamba feature dimension is reduced to 128, these reductions increase to approximately 49% for both training and inference memory usage. Besides, the inference time of OccMamba-128 is approximately half of the M-CONet and OccMamba-128 remarkably surpasses M-CONet in performance.

### v2

@@ 3.3 OccMamba encoder with height-prioritized reordering
Semantic occupancy prediction is challenging due to the high dimensionality and density of voxel grids, often involving millions of voxels. Previous methods [14, 28] resort to projection techniques but suffer from information loss. In contrast, Mamba’s linear computational complexity enables direct processing of large voxel features, avoiding the limitations of deformable attention and convolution, such as the need for key point selection or limited receptive fields.
@@
Height-prioritized reordering scheme. Before feeding 3D voxel features into the Mamba block, it’s necessary to reorder them into 1D sequences. A poor reordering strategy can disrupt the intrinsic spatial relationships between adjacent voxels in the 3D space, especially when dealing with a large number of voxel grids, thereby impacting Mamba’s performance. Inspired by the Hilbert curve [12], we propose a height-prioritized 2D Hilbert expansion, tailored to the flat spatial structure typical of semantic occupancy prediction tasks, where height information provides valuable categorical clues, as it often correlates with object categories, reveals terrain features, and distinguishes different spatial regions in the scene. Specifically, we divide the ”xyz” coordinates into the ”xy” plane and a ”z” dimension. Starting on the xy-plane at z=0, the process extends vertically along the z-axis, creating vertical sequences. These sequences are then ordered and interconnected following the 2D Hilbert curve on the xy plane, forming the height-prioritized 2D Hilbert expansion, as illustrated in Fig. 3(d). For comparison, Fig. 3(a)-(c) illustrate the results of reordering according to the ’xyz’ sequence, the ’zxy’ sequence, and the 3D Hilbert curve, respectively. Our reordering strategy prioritizes z-axis spatial information while maintaining strong spatial proximity in the xy plane. By this way, we can better leverage Mamba’s contextual modeling.
As a result, with the spatial size $W$, $H$, $D$ and 1D data length $L$, the input $\mathbf{V}\in\mathbb{R}^{B\times W\times H\times D\times C}$ and the output $\mathbf{V^{\prime}}\in\mathbb{R}^{B\times L\times C}$ in the above reordering scheme $\mathcal{R}_{3D\rightarrow 1D}$ and $\mathcal{R}_{1D\rightarrow 3D}$ can be expressed as:
$\mathbf{V}=\mathcal{R}_{1D\rightarrow 3D}(\mathbf{V^{\prime}})=\mathcal{R}_{1D\rightarrow 3D}(\mathcal{R}_{3D\rightarrow 1D}(\mathbf{V}))$ (7)
@@
Hierarchical mamba module. To effectively utilize the contextual information across spatial resolutions, we introduce the hierarchical Mamba module, which includes an encoder and decoder. The encoder comprises multiple groups, each with two Mamba blocks and a downsampling operation between groups for multi-scale voxel features. Correspondingly, the Mamba decoder mirrors this structure, utilizing paired Mamba blocks and upsampling operations. The upsampling incorporates skip connection from the corresponding layer in the Mamba encoder, ensuring consistency in feature dimensions across scales. Crucially, the reordering scheme is applied both before and after each Mamba blocks group, maintaining voxel representation during both downsampling and upsampling phases.
As to a single Mamba block, it consists of LayerNorm(LN), linear layers, 1D convolution, Silu activation, Selective SSM, and residual connections. Given an input $\mathbf{V}\in\mathbb{R}^{B\times L\times C}$, where $B$, $L$ and $C$ denote the batch size, the length of 1D data, and the feature dimension, respectively, the output $\mathbf{V^{\prime}}\in\mathbb{R}^{B\times L\times C}$ can be computed as below:
@@
$\mathbf{V^{\prime}}$ $=\text{Linear}(\mathbf{V^{\prime}_{2}}\odot\text{Silu}(\text{Linear}(\mathbf{V^{\prime}_{1}}))),$ (10)
where $\odot$ represents the dot product. For convenience, we denote Eq. (8)-Eq. (10) as $\mathcal{M}$. Then, Given the input $\mathbf{V}\in\mathbb{R}^{B\times W\times H\times D\times C}$ and output $\mathbf{V^{\prime}}\in\mathbb{R}^{B\times W\times H\times D\times C}$, the computation for a single Mamba block with our reordering scheme $\mathcal{R}_{3D\rightarrow 1D}$ and $\mathcal{R}_{1D\rightarrow 3D}$ can be represented as
$\mathbf{V^{\prime}}=\mathcal{R}_{1D\rightarrow 3D}\big{(}\mathcal{M}(\mathcal{R}_{3D\rightarrow 1D}(\mathbf{V}))\big{)}$ (11)
Local context processor. Although the hierarchical Mamba module has improved the network’s ability to process multi-scale global information, handling more local information can further enhance the accuracy of semantic occupancy prediction. To this end, we design a lightweight local context processor. It divides the Mamba features $\mathbf{V}_{\mathcal{M}}\in\mathcal{R}^{B\times W\times H\times D\times C}$ along the XY plane into patches $\mathcal{V}=\{\mathcal{V}_{s_{i},w_{i}}|s_{i}\in S,w_{i}\in W\}$, using a list of window sizes $W=\{w_{i}\}$ and a corresponding list of sliding strides $S=\{s_{i}\}$ to scan the plane. For each item $\mathbf{V}_{p,q}\in\mathcal{V}_{s_{i},w_{i}}$, we have the following equations:
$\mathbf{V}_{p,q}$ $=\mathbf{V}_{\mathcal{M}}[:,p\cdot s_{i}:p\cdot s_{i}+w_{i},q\cdot s_{i}:q\cdot s_{i}+w_{i},:,:],$
$p$ $=0,1,...\left\lfloor\frac{W-w_{i}}{s_{i}}\right\rfloor,q=0,1,...\left\lfloor\frac{H-w_{i}}{s_{i}}\right\rfloor.$ (12)
Next, we reorder $\mathbf{V}_{p,q}\in\mathcal{R}^{B\times w_{i}\times w_{i}\times D\times C}$ to $\mathcal{R}^{B\times L\times C}$, and then stack $\mathcal{V}_{s_{i},w_{i}}$ to $\mathcal{R}^{(B\times N)\times L\times C}$. Each $\mathcal{V}_{s_{i},w_{i}}$ is processed by the dedicated two-layer Mamba blocks, and the outputs are reshaped to $\mathcal{V^{\prime}}_{s_{i},w_{i}}=\{\mathbf{V^{\prime}}_{p,q}\in\mathcal{R}^{B\times w_{i}\times w_{i}\times D\times C}\}$. Finally, $\mathcal{V^{\prime}}=\{\mathcal{V^{\prime}}_{s_{i},w_{i}}\}$ patches are reassembled into the original spatial size and concatenated along the channel dimension to produce $\mathbf{V^{\prime}}_{\mathcal{L}}\in\mathcal{R}^{B\times W\times H\times D\times C^{\prime}}$. A 3D convolution with a kernel size of $1\times 1\times 1$ is then applied along the channel dimension to reduce it, resulting in $\mathbf{V}_{\mathcal{L}}\in\mathcal{R}^{B\times W\times H\times D\times C}$.
$\mathcal{V^{\prime}}_{s_{i},w_{i}}$ $=\text{Reshape}(\mathcal{M}(\text{Stack}(\text{Reorder}(\mathcal{V}_{s_{i},w_{i}}))))$ (13)
[…]
Endowed with the hierarchical Mamba module and local context processor, our OccMamba utilizes both global and local information from dense scene grids, thereby capturing broad context while preserving local details. This integrated approach enhances spatial understanding and mitigates potential performance degradation with large data volumes, by effectively balancing global awareness with local precision.
MonoScene [4] C 18.4 6.9 7.1 3.9 9.3 7.2 5.6 3.0 5.9 4.4 4.9 4.2 14.9 6.3 7.9 7.4 10.0 7.6
TPVFormer [14] C 15.3 7.8 9.3 4.1 11.3 10.1 5.2 4.3 5.9 5.3 6.8 6.5 13.6 9.0 8.3 8.0 9.2 8.2
SparseOcc [39] C 21.8 14.1 16.1 9.3 15.1 18.6 7.3 9.4 11.2 9.4 7.2 13.0 31.8 21.7 20.7 18.8 6.1 10.6
3DSketch [5] C&D 25.6 10.7 12.0 5.1 10.7 12.4 6.5 4.0 5.0 6.3 8.0 7.2 21.8 14.8 13.0 11.8 12.0 21.2
AICNet [15] C&D 23.8 10.6 11.5 4.0 11.8 12.3 5.1 3.8 6.2 6.0 8.2 7.5 24.1 13.0 12.8 11.5 11.6 20.2
LMSCNet [35] L 27.3 11.5 12.4 4.2 12.8 12.1 6.2 4.7 6.2 6.3 8.8 7.2 24.2 12.3 16.6 14.1 13.9 22.2
JS3C-Net [49] L 30.2 12.5 14.2 3.4 13.6 12.0 7.2 4.3 7.3 6.8 9.2 9.1 27.9 15.3 14.9 16.2 14.0 24.9
M-CONet [45] C&L 29.5 20.1 23.3 13.3 21.2 24.3 15.3 15.9 18.0 13.3 15.3 20.7 33.2 21.0 22.5 21.5 19.6 23.2
Co-Occ [31] C&L 30.6 21.9 26.5 16.8 22.3 27.0 10.1 20.9 20.7 14.5 16.4 21.6 36.9 23.5 5.5 23.7 20.5 23.5
OccMamba-128 (ours) C&L 34.7 25.2 29.1 19.1 25.5 28.5 18.1 24.7 23.4 19.8 19.3 24.5 37.0 25.4 25.4 25.4 28.1 29.9
OccMamba-384 (ours) C&L 35.7 26.2 30.2 20.5 26.5 29.5 18.8 26.0 23.7 19.9 20.6 25.4 38.4 26.5 27.0 26.6 28.9 30.5
Table 1: Quantitative comparisons on OpenOccupancy validation set with v0.0 annotations. C, D, L denote camera, depth and LiDAR, respectively. OccMamba-384 means OccMamba with mamba feature dimension being 384. The best and second-best are in bold and underlined, respectively.
Method Input Modality mIoU
MonoScene [4] C 11.1
SurroundOcc [46] C 11.9
OccFormer [52] C 12.3
RenderOcc [32] C 12.8
LMSCNet [35] L 17.0
JS3C-Net [49] L 23.8
SSC-RS [29] L 24.2
Co-Occ [31] C&L 24.4
M-CONet [45] C&L 20.4
OccMamba-128 (ours) C&L 24.6
Method Input Modality mIoU
SSCNet [38] L 15.2
LMSCNet [35] L 16.5
MotionSC [47] L 17.6
JS3C-Net [49] L 22.7
OccMamba-128 (ours) L 23.4
@@ 3.4 Training objective
Our ultimate training objective is composed of five terms, including the cross-entropy loss $\mathcal{L}_{\text{CE}}$, the lovasz-softmax loss $\mathcal{L}_{\text{iou}}$ [2], the geometric and semantic affinity loss $\mathcal{L}_{\text{geo}}$ and $\mathcal{L}_{\text{sem}}$ [4], and the depth supervision loss $\mathcal{L}_{\text{depth}}$ [19]. These terms are crucial for optimizing our model’s performance in various aspects: $\mathcal{L}_{\text{CE}}$ ensures accurate classification, $\mathcal{L}_{\text{iou}}$ enhances semantic segmentation, $\mathcal{L}_{\text{geo}}$ improves spatial alignment, $\mathcal{L}_{\text{sem}}$ refines semantic understanding, and $\mathcal{L}_{\text{depth}}$ guides depth estimation and spatial relationships. The ultimate training objective is presented as follows:
@@ 4.1 Experimental setup
Benchmarks. We conduct experiments on OpenOccupancy [45], SemanticKITTI [1] and SemanticPOSS [33]. OpenOccupancy is built upon the nuScenes dataset [3], inheriting the data format of nuScenes. It comprises 700 training sequences and 150 validation sequences, with annotations for 17 classes. The occupancy annotations are represented in a 512$\times$512$\times$40 voxel grid, with each voxel sized at 0.2 meters. Notably, each frame in OpenOccupancy has a data range four times larger than the other datasets, imposing a heavier computational burden. In SemanticKITTI, sequences 00-10 (excluding 08), 08, and 11-21 are allocated for training, validation, and testing, respectively. For the occupancy annotations, a 256$\times$256$\times$32 grid is used, with voxels measuring 0.2 meters each. After pre-processing, a total of 19 classes are utilized for training and evaluation. SemanticPOSS, which closely resembles SemanticKITTI, employs sequences 00-05 and 02 as training and validation sets, respectively, with annotations for 11 classes.
@@
Implementation details. We use ResNet-50 [11] as the image backbone. Both the Mamba Encoder and Mamba Decoder in the hierarchical mamba module consist of four groups and each group contains two Mamba blocks. The local context processor utilizes window sizes of [3, 5, 7], with two Mamba blocks implemented. For the OpenOccupancy dataset, we maintain identical setting to M-CONet [45]: for each frame, we utilize six surround view camera images as input, coupled with a fusion of ten frames of LiDAR points spanning the range of [-51.2m, 51.2m] along the X and Y axes, and [-2.0m, 6.0m] along the Z axis. Within the Mamba block, experiments are conducted with mamba feature dimension set to 128 and 384, respectively. On the SemanticKITTI dataset, our inputs consist of forward-facing stereo camera images alongside a single-frame LiDAR points, with their spatial extents defined as [0m, -25.6m, -2m, 51.2m, 25.6m, 4.4m]. We set the mamba feature dimension to 128 and employ Test Time Augmentation (TTA) to boost the performance. The used augmentations include the flipping of the X and Y axes, and four augmented inputs (including the original input) are sent to the model. Finally, for the SemanticPOSS, we exclusively leverage single-frame LiDAR points like SemanticKITTI.
4.2 Experimental results
OpenOccupancy. As shown in Table 1, we compare OccMamba on OpenOccupancy with v0.0 annotation to previous methods which use the settings of camera-only [4, 14, 5, 15], LiDAR-only [35, 49], and multi-modal fusion [45, 31]. Our OccMamba achieves the best performance. Particularly, OccMamba-384 outperforms Co-Occ [31] by 5.1% in IoU and 4.3% in mIoU, respectively. Additionally, OccMamba-128 also surpasses Co-Occ by 4.1% in IoU and 3.3% in mIoU. These results underscore the efficacy of our approach in handling large-scale scenarios.
SemanticKITTI & SemanticPOSS. As evident from Table 2, our OccMamba-128 still outperforms the second-best method, i.e., Co-Occ [31], by 0.2% mIoU in SemanticKITTI test set. Note that we do not compare our OccMamba with SCPNet [48] and Occfiner [37] as they employ tricks such as label rectification, knowledge distillation or multi-frame concatenation. Strong performance is also observed in SemanticPOSS, as shown in Table 3. These results sufficiently demonstrate the superiority of our approach. The detailed classwise performance is put in the supplementary material.
Visual comparisons. From Fig. 4, our OccMamba-384 produces more accurate predictions than M-CONet [45] on OpenOccupancy validation set. Our OccMamba not only provides a more detailed representation of object shapes and semantics, but also performs preciser predictions on occluded objects and surfaces.
Reordering Schemes mIoU
XYZ sequence 24.5
ZXY sequence 24.7
3D Hilbert 24.8
Height-prioritized 2D Hilbert expansion 25.2
Table 4: Ablation study on different reordering schemes.
Method CNN Hierarchical Mamba Module Local Context Processor mIoU
M-CONet [45] 20.1
OccMamba-128 25.0
OccMamba-128 25.2
OccMamba-384 26.2
Table 5: Ablation study on OccMamba Encoder.
@@ 4.3 Ablation study
The reported results are on the OpenOccupancy validation set with v0.0 annotations unless otherwise specified.
Reordering schemes. To accelerate the training process, we set the mamba feature dimension to 128 and then compare our method with three alternative methods: ordering by XYZ sequence, ordering by ZXY sequence, and ordering using the 3D Hilbert curve. As shown in Table 4, our method achieves the highest mIoU of 25.2%. Additionally, it is apparent that reordering schemes that prioritize the z-axis outperform the other methods. This could be attributed to the fact that in semantic occupancy prediction, the overall shape of the scenes typically manifests as a flattened cuboid. Therefore, reordering schemes that prioritize the z-axis allow Mamba to better capture spatial proximity information, leading to improved prediction performance.
OccMamba encoder. Consistent with previous settings, we conduct ablation studies on mamba feature dimension and two modules within our OccMamba encoder. As shown in Table 5, by replacing the CNN-based occupancy encoder from M-CONet [45] with our hierarchical Mamba module using a mamba feature dimension of 128, the mIoU improves from 20.1% to 25.0%. The incorporation of our local context processor further enhances the performance, achieving 25.2% mIoU. When the Mamba feature dimension is increased from 128 to 384, the mIoU improves further, reaching 26.2%.
Generalization. We apply our hierarchical Mamba module to the MonoScene[4] framework. Specifically, we replace the 3D voxel decoder in this framework with our implementation, using a mamba feature dimension of 128, and train on 25% of the SemanticKITTI training set with the training strategy unchanged. As shown in Table 6, our OccMamba encoder can improve the performance of MonoScene from 11.1% mIoU to 11.9% mIoU on SemanticKITTI val set.
Memory usage and inference time. In Table 7, we present a comparative analysis of memory usage and inference time between OccMamba and M-CONet during both training and inference phases. The networks are trained on 8 A40 GPUs and perform inference on a single RTX 4090 GPU.When the mamba feature dimension is set to 384, OccMamba shows similar memory usage to M-CONet during training but reduces inference memory usage by about 24%. When the mamba feature dimension is lowered to 128, the memory savings are more pronounced, with a reduction of approximately 38% during training and 44% during inference compared to M-CONet. Additionally, OccMamba-128 significantly improves inference speed, taking about 35% of the time required by M-CONet, which highlights its efficiency in resource-constrained scenarios. In terms of performance, OccMamba-384 achieves the highest mIoU, while OccMamba-128 configurations also surpass M-CONet in performance, providing a balance between efficiency and accuracy.
Method CNN Hierarchical Mamba Module Local Context Processor mIoU
MonoScene [4] 11.1
MonoScene [4] 11.7
MonoScene [4] 11.9
@@
M-CONet [45] 37.3 12.1 416.0 20.1
OccMamba-384 37.7 9.2 448.3 26.2
OccMamba-128 23.1 6.8 269.3 25.2
Table 7: Comparison of memory usage and inference time.